OpenStudy (anonymous):

represents the path of an airplane that passes through the coordinates (2, 1) and (4, 5). What is the slope of a line that represents the path of another airplane that is traveling parallel to the first airplane?

OpenStudy (texaschic101):

parallel lines have the same slope. So we will find the slope by using the slope formula. slope = (y2 - y1) / (x2 - x1) (2,1)...x1 = 2 and y1 = 1 (4,5)...x2 = 4 and y2 = 5 now we sub slope = (5 - 1) / (4 - 2) slope = 4/2 -- reduce slope = 2

OpenStudy (texaschic101):

any questions ?